### v1

# Exponentially accelerated relaxation and quantum Mpemba effect in open quantum systems

Abstract
We investigate the quantum Mpemba effect in the relaxation of open quantum systems whose effective dynamics is described by Davies maps. We present a class of unitary transformations based on permutation matrices that, acting on the initial state of the system, (i) suppress the contribution of slowest decaying modes of the nonunitary dynamics; (ii) ensure that it is as distinguishable as possible from the steady state. The first requirement guarantees an exponentially accelerating convergence to the steady state, while the second implies that a quantum system initially farther from equilibrium approaches it more rapidly than an initial state closer to it. This protocol provides a genuine Mpemba effect, and its numerical simulation requires low computational effort. We prove that, for any initial state, there always exists a permutation matrix that maximizes its distance from the equilibrium for a given information-theoretic distinguishability measure. We illustrate our findings for the nonunitary dynamics of the transverse field Ising chain and XXZ chain, each weakly coupled to a bosonic thermal bath, showing the quantum Mpemba effect captured by the Hilbert-Schmidt distance, quantum relative entropy, and trace distance. Our results provide a universal and versatile framework to engineer the genuine quantum Mpemba effect in open quantum systems.
I Introduction
Intuition suggests that the cooling of a physical system is a monotonic process: the closer to the equilibrium, the faster it should reach it. However, observations dating back to Aristotle over 2000 years ago Aristotle (1952), as well as studies developed by Mpemba and Osborne Mpemba and Osborne (1969), and Kell Kell (1969) in 1969, reveal counterintuitive behavior: under certain conditions, initially hotter systems can cool more rapidly than initially colder systems. This effect is called as the Mpemba effect. Since then, the phenomenon has been reported in different physical contexts, including crystalline polymers, manganites, and colloidal suspensions Hu et al. (2018); Chaddah et al. ; Kumar and Bechhoefer (2020). Despite its widespread occurrence, the underlying mechanism remains the subject of intense debate Burridge and Linden (2016); Bechhoefer et al. (2021). Recent advances in stochastic thermodynamics suggest that initially hotter systems can exploit dynamical shortcuts, accessing regions of state space that allow them to relax more rapidly than colder systems Lu and Raz (2017); Klich et al. (2019). We also mention the interplay between Mpemba and Kovacs effects in the so-called time-delayed cooling law Santos (2024).
Figure 1: (Color online) Overview of the quantum Mpemba effect (QME). (a) We consider a probe state $\rho({t_{0}})$ that is transformed into ${\rho^{\prime}}({t_{0}})=U\rho({t_{0}}){U^{\dagger}}$ by means of the unitary matrix $U={U_{1}}{P_{\pi}}{\Lambda^{\dagger}}$. This state contributes to eliminate the slowest decaying mode in Davies maps with generator $\mathcal{L}[\bullet]$. Therefore, the relaxation to equilibrium is accelerated exponentially, controlled by the real part of eigenvalue $\lambda_{4}$. The unitary $U$ maximizes the distance between ${\rho^{\prime}}({t_{0}})$ and the steady state $\texttt{R}_{1}$}, such that $\mathcal{D}({\rho^{\prime}}({t_{0}}),{\texttt{R}_{1}})>\mathcal{D}({\rho}({t_{0}}),{\texttt{R}_{1}})$. QME occurs if there exist a time $t_{\text{QME}}$} such that for all $t>{t_{\text{QME}}}$, one finds $\mathcal{D}({e^{t\mathcal{L}}}[{\rho^{\prime}}({t_{0}})],{\texttt{R}_{1}})<\mathcal{D}({e^{t\mathcal{L}}}[\rho({t_{0}})],{\texttt{R}_{1}})$. (b) The protocol to accelerate convergence to equilibrium involves the following steps: (1) unitary matrix $\Lambda$ maps the probe state to its diagonal form; (2) permutation matrix $P_{\pi}$} rearranges the diagonal entries of ${\Lambda^{\dagger}}\rho({t_{0}}){\Lambda}$, so that the transformed state is as far away from equilibrium as possible; (3) the resulting state from the previous step is rotated by the unitary matrix $U_{1}$ to the eigenbasis of the Hamiltonian.
The Mpemba effect has found an analogue in the quantum regime Nava and Fabrizio (2019). The so-called quantum Mpemba effect (QME) is defined as a physical process where a system initially further from equilibrium relaxes faster than a system closer to equilibrium. In Markovian open quantum systems, by acting a suitable unitary operator on the pure initial state of the system, one verifies the suppression of the slowest decay mode of the nonunitary dynamics Carollo et al. (2021); Kochsiek et al. (2022, 2022); Bao and Hou (2025). This mechanism ensures an exponentially faster convergence to the stationary state, which defines the QME. It is worth noting, however, that the slowest decaying mode does not always contribute to the relaxation mechanism. This was observed, for example, in the effective Markovian dynamics of a single-level quantum dot coupled to two reservoirs Chatterjee et al. (2023). We refer to Refs. Ares et al. (2025a); Teza et al. for recent reviews on QME.
It turns out that eliminating slowest decaying mode is not enough to ensure the Mpemba effect. This issue motivated the proposal of a genuine quantum Mpemba effect, which involves accelerated relaxation to equilibrium, and crossover between relaxation curves related to useful figures of merit that capture the distance from the probe state to the steady state Moroder et al. (2024). This setting involves mixed initial states and Davies maps whose spectral gap is dictated by a complex pair of eigenvalues related to a non-Hermitian jump operator. We highlight a recent study addressing QME for open nonequilibrium quantum systems coupled to two different reservoirs Nava and Egger (2024); Wang and Wang (2024), which paved the way for investigations concerning the so-called Pontus-Mpemba effect Nava and Egger (2025); Nava et al. .
The quantum Mpemba effect also manifests itself in the restoration of broken symmetries in many-body quantum systems. In this setting, QME occurs when the symmetry is locally restored faster for the initial state the breaks it more Rylands et al. (2024); Ares et al. (2023); Liu et al. (2024); Yamashika et al. (2024); Ares et al. (2025b); Yamashika et al. (2025); Russotto et al. (2025). We also mention recent works addressing QME in the study of quantum complexity Aditya et al. ; Mackinnon and Paternostro , local relaxation of closed quantum systems Ares et al. (2025c), and quantum speed limits Bettmann and Goold (2025). Experimental evidence of the quantum Mpemba effect was reported in trapped-ion quantum simulator Joshi et al. (2024); Zhang et al. (2025), including its inverse version Aharony Shapira et al. (2024), and more recently QME was addressed in a Nuclear Magnetic Resonance platform Schnepper et al. .
In this work, we present a framework that allows for exponentially faster relaxation to equilibrium and QME in open quantum systems described by Davies maps [see Fig. 1(a)]. Our approach takes into account a unitary operator based on permutation matrices which, once applied to the initial state of the system, rearranges its spectrum and induces the suppression of the slowest decaying mode of the dynamics [see Fig. 1(b)]. On the one hand, we realize that exponential speed up is achieved regardless the permutation matrix. On the other hand, we prove that there always exists a permutation matrix that maximizes the distance from the initial state to equilibrium. This distance is characterized by paradigmatic distinguishability measures, namely, Hilbert-Schmidt distance, quantum relative entropy, and trace distance. Together, these features guarantee the occurrence of the genuine quantum Mpemba effect. We present numerical simulations to support our findings. We emphasize that our approach can also be appropriately modified to encompass other class of nonunitary evolutions, thus recovering previous results from the literature. Finally, we show that, in order to achieve an exponentially rapid evolution towards equilibrium, an initial state does not necessarily need to be incoherent respective to the eigenbasis of the Hamiltonian.
The outline of the paper is as follows. In Sec. II, we revise the general theory of Davies maps, particularly addressing the spectral properties of the Liouvillian that governs the effective nounitary dynamics of the quantum system. In Sec. III, we present a framework that engineers an exponential speed up relaxation by eliminating the slowest decay modes of the Liouvillian. In Sec. IV, we show that there is always possible to find a permutation matrix such that we have a dressed state far from the initial state. Hence, once our protocol induces quicker relaxation towards equilibrium, these two features trigger genuine quantum Mpemba effect. In Sec. V, we illustrate our findings for the transverse field Ising model, and the XXZ model, each of these spin chains weakly coupled to a thermal bath. In Sec. VI, we showed that there are initial states for which an exponentially rapid relaxation to the equilibrium is already observed, but which are not incoherent with respect to the energy eigenbasis. Finally, in Sec. VII we summarize our conclusions.
II Markovian quantum dynamics
Here we consider the effective nonunitary dynamics described by Davies maps, that is, quantum dynamical semigroups that describe the relaxation to the equilibrium of a given $d$-dimensional quantum system weakly coupled to a thermal bath at temperature $T$ Davies (1979); Breuer and Petruccione (2007). The nonunitary dynamics is governed by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation $d\rho(t)/dt=\mathcal{L}[\rho(t)]$, with the Liouvillian operator given by
$\mathcal{L}[\bullet]=-i[H,\bullet]+{\sum_{s=1}^{2}}\,{\sum_{\begin{subarray}{c}n,m=1\\ m<n\end{subarray}}^{d}}\,{\texttt{D}_{nm}^{(s)}}[\bullet]\penalty 10000\ ,$ (1)
where the $s$-th dissipator operator is written as follows
${\texttt{D}_{nm}^{(s)}}[\bullet]={L_{nm}^{(s)}}\bullet{L_{nm}^{(s)\dagger}}-\frac{1}{2}\{{L_{nm}^{(s)\dagger}}{L_{nm}^{(s)}},\bullet\}\penalty 10000\ ,$ (2)
with $m,n\in\{1,\ldots,d\}$, $\bullet$$\bullet$] and $\{\bullet,\bullet\}$ define the commutator and anticommutator, respectively, and we set $\hbar=1$ throughout the manuscript. Here, $H={\sum_{l=1}^{d}}\,{\varepsilon_{l}}|{\psi_{l}}\rangle\langle{\psi_{l}}|$ is the Hamiltonian of the system, where $\{{\varepsilon_{l}}\}_{l=1,\ldots,d}$} are the energies, and $\{|{\psi_{l}}\rangle\}_{l=1,\ldots,d}$} is the complete set of eigenvectors. In turn, the jump operators are given by Moroder et al. (2024)
${L_{nm}^{(1)}}=\sqrt{\xi_{nm}}\,|{\psi_{n}}\rangle\langle{\psi_{m}}|\penalty 10000\ ,\quad{L_{nm}^{(2)}}=\sqrt{\chi_{nm}}\,|{\psi_{m}}\rangle\langle{\psi_{n}}|\penalty 10000\ ,$ (3)
for $m,n=\{1,\ldots,d\}$, with $m<n$, and
${\chi_{nm}}=\frac{\gamma}{{e^{({\varepsilon_{m}}-{\varepsilon_{n}})/{k_{B}}T}}\pm 1}\penalty 10000\ ,\quad{\xi_{nm}}=\gamma\mp{\chi_{nm}}\penalty 10000\ ,$ (4)
where $\gamma$ is the strength of the coupling between system and environment, $k_{B}$} is the Boltzmann constant, while the signs $\pm$ refer to the Fermi-Dirac (++) and Bose-Einstein (-) distribution functions. Importantly, Davies maps fulfill the quantum detailed balance, while both coherent and dissipative parts of the generator commute Spohn (1978); Breuer (2003).
The formal solution of the Markovian master equation is given by $\rho(t)={e^{t\mathcal{L}}}[\rho({t_{0}})]$, where $\rho({t_{0}})$ is the initial state of the system at time $t_{0}$. This instantaneous state can be recast in terms of the spectral decomposition of the Lindblad operator as follows
$\rho(t)$ $={\texttt{R}_{1}}+{e^{t{\lambda_{2}}}}\text{Tr}({\texttt{L}_{2}}\rho({t_{0}})){\texttt{R}_{2}}+{e^{t{\lambda^{*}_{2}}}}\text{Tr}({\texttt{L}_{2}^{\dagger}}\rho({t_{0}})){\texttt{R}_{2}^{\dagger}}$
$+{\sum_{s=4}^{{d^{2}}}}\,{e^{t{\lambda_{s}}}}\text{Tr}({\texttt{L}_{s}}\rho({t_{0}})){\texttt{R}_{s}}\penalty 10000\ ,$ (5)
where $d$ is the dimension of the Hilbert space of the system, while $\texttt{R}_{s}$} and $\texttt{L}_{s}$} denote, respectively, the $s$-th right and left eigenmatrices of the Liouvillian in Eq. (1) that are related to the complex eigenvalue $\lambda_{s}$, thus satisfying $\mathcal{L}[{\texttt{R}_{s}}]={\lambda_{s}}{\texttt{R}_{s}}$ and ${\mathcal{L}^{\dagger}}[{\texttt{L}_{s}}]={\lambda_{s}}{\texttt{L}_{s}}$. We note that $\mathcal{L}^{\dagger}$$\bullet$] defines the adjoint map. The set of right and left eigenmatrices span a basis for the space of matrices, satisfying the orthogonality constraint $\text{Tr}({\texttt{L}_{s}^{\dagger}}{\texttt{R}_{l}})={\delta_{sl}}$ for all $s,l\in\{1,\ldots,{d^{2}}\}$.
To generate a completely positive dynamics, it can be proven that the Lindblad superoperator must exhibit complex eigenvalues with negative real parts, i.e., $\text{Re}({\lambda_{k}})\leq 0$, for all $k=\{1,\ldots,{d^{2}}\}$. However, to ensure that such a nonunitary map is trace-preserving, i.e., $\text{Tr}[\rho(t)]=1$ for all $t\geq 0$, it is verified that at least one of these eigenvalues must be zero, say ${\lambda_{1}}=0$. We assume that this eigenvalue is nondegenerate, which implies an unique steady state ${\rho_{\infty}}:={\lim_{t\rightarrow\infty}}\rho(t)$ given exactly by the eigenmatrix $\texttt{R}_{1}$ [see Eq. (II)], with $\text{Tr}({\texttt{R}_{1}})=1$, also implying that the left eigenmatrix is the identity operator, ${\texttt{L}_{1}}=\mathds{1}$ Prosen (2008). For our purposes, the set of eigenvalues of the Lindblad superoperator is arranged in ascending order respective to the absolute value of their real part such that $0={\lambda_{1}}<|\text{Re}({\lambda_{2}})|\leq|\text{Re}({\lambda_{3}})|\leq\ldots$. The mode with eigenvalue $\lambda_{2}$} [see Eq. (II)], having the smallest nonzero real part, sets the longest relaxation time, given by ${t_{2}}=1/{|\text{Re}({\lambda_{2}})|}$.
The slowest decay mode controls the relaxation timescale towards to the steady state. For long times, the dynamics is expected to be dominated by such a mode, unless its overlap with the initial state becomes negligible. In fact, an exponential speed up in relaxation towards the equilibrium state is expected whenever the constraints $\text{Tr}({\texttt{L}_{2}^{\dagger}}\rho({t_{0}}))=0$ and $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))=0$ are satisfied Carollo et al. (2021). More generally, this speed up can be engineered by a unitary matrix $U$ such that the transformed state ${\rho^{\prime}}({t_{0}})=U\rho({t_{0}}){U^{\dagger}}$ has a vanishing overlap with the lowest decaying eigenmatrix $\texttt{L}_{2}$}, namely,
$\text{Tr}({\texttt{L}_{2}^{\dagger}}U\rho({t_{0}}){U^{\dagger}})=0\penalty 10000\ ,$ (6)
and
$\text{Tr}({\texttt{L}_{2}}U\rho({t_{0}}){U^{\dagger}})=0\penalty 10000\ .$ (7)
This guarantees that the slowest mode is effectively suppressed, thus accelerating relaxation to the equilibrium Kochsiek et al. (2022). Next, we will address the issue of accelerating the relaxation process. Here, we focus on open quantum systems described by Davies maps, but we emphasize that our approach can encompass other types of dynamical maps [see Appendix A].
III Speeding up the relaxation process
We now focus on investigating a protocol to accelerate the relaxation of the open quantum system towards equilibrium. For our purposes, we recast the Hamiltonian as $H={U_{1}}\varepsilon{U_{1}^{\dagger}}$, where $\varepsilon=\text{diag}({\varepsilon_{1}},\ldots,{\varepsilon_{d}})$ is the diagonal matrix that contains its energies, with ${\varepsilon_{l}}>{\varepsilon_{l+1}}$ for all $l=\{1,\ldots,d-1\}$, while $U_{1}$} is the unitary matrix formed by the respective eigenvectors. Overall, for a given initial state, eliminating the slowest decaying mode requires preparing a state ${\rho^{\prime}}({t_{0}})={U}\rho({t_{0}}){U^{\dagger}}$ that is orthogonal to the eigenmatrix $\texttt{L}_{2}$} of the Liouvillean, for a given unitary matrix $U$. To achieve the results in Eqs. (6) and (7), we consider the unitary operator
$U={U_{1}}{P_{\pi}}{\Lambda^{\dagger}}\penalty 10000\ ,$ (8)
where $P_{\pi}$} is a given permutation matrix, and $\Lambda$ is the unitary matrix formed by the eigenstates of $\rho({t_{0}})$, such that $\rho({t_{0}})={\Lambda}{D}{\Lambda^{\dagger}}$, with $D=\text{diag}({\lambda_{1}},\ldots,{\lambda_{d}})$ is a diagonal matrix whose elements are the eigenvalues of the probe state. By applying this unitary to the probe state, we obtain the density matrix
${\rho^{\prime}}({t_{0}})$ $={U}\rho({t_{0}}){U^{\dagger}}$
$={U_{1}}{P_{\pi}}{\Lambda^{\dagger}}\rho({t_{0}})\Lambda{P_{\pi}^{\dagger}}{U_{1}^{\dagger}}$
$={U_{1}}{P_{\pi}}{D}{P_{\pi}^{\dagger}}{U_{1}^{\dagger}}\penalty 10000\ ,$ (9)
which from now on we refer to dressed initial state. In Fig. 1(b), we illustrate the role of the unitary matrix $U$. In Appendix A, we show that Eq. (8) can be reformulated to recover the results discussed in Ref. Carollo et al. (2021).
The unitary transformation in Eq. (III) first maps $\rho({t_{0}})$ to its diagonal form. The permutation matrix $P_{\pi}$} then rearranges the order of the diagonal entries of $D$, which in turn is mapped onto the diagonal matrix ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},\ldots,{\lambda_{\pi(d)}})$. The permutation matrix $P_{\pi}$} is constructed in order to maximize the distinguishability between ${\rho^{\prime}}({t_{0}})$ and the steady state $\texttt{R}_{1}$. This means that the diagonal entries of $D$ will be rearranged in ascending order, ${\lambda_{\pi(1)}}<\ldots<{\lambda_{\pi(d)}}$, or even descending order, ${\lambda_{\pi(1)}}>\ldots>{\lambda_{\pi(d)}}$, depending on the figure of merit used to evaluate the distance from the transformed state to equilibrium. It will become clear later that $P_{\pi}$} plays an essencial role in ensuring that the genuine the Mpemba effect occurs for a given nonunitary dynamics [see Sec. IV]. Finally, the resulting matrix is rotated by $U_{1}$ to the quantum state in Eq. (III), which is incoherent with respect to the eigenbasis of the Hamiltonian $H$. In comparison with the results from Ref. Moroder et al. (2024), we note that the numerical simulation of our protocol requires low computational effort. The calculations involve prior knowledge of the spectrum of operators $H$ and $\rho({t_{0}})$, while the permutation matrix is determined by the appropriate ordering of the eigenvalues of $\rho({t_{0}})$ to maximize the distance from ${\rho^{\prime}}({t_{0}})$ to equilibrium.
The unitary in Eq. (8) causes the slowest decaying mode of the Liouvillian to be mapped onto the matrix ${\texttt{L}^{\prime}_{2}}={U_{1}^{\dagger}}{\texttt{L}_{2}}{U_{1}}$, where $\texttt{L}^{\prime}_{2}$} is an upper/lower triangular matrix that has all elements zero except for one of its off-diagonal entries. In detail, we have that ${\texttt{L}^{\prime}_{2}}=|{j_{0}}\rangle\langle{l_{0}}|$, for a given pair of states $|{j_{0}}\rangle$ and $|{l_{0}}\rangle$, with ${j_{0}}\neq{l_{0}}$, where $\{|{j}\rangle\}_{j=1,\ldots,d}$} defines the computational basis related to the $d$-dimensional Hilbert space of the system [e.g., the set of eigenstates of the observable ${S_{z}}=(1/2){\sum_{j=1}^{d}}{\sigma_{j}^{z}}$]. In this setting, given that ${U}\rho({t_{0}}){U^{\dagger}}={U_{1}}{P_{\pi}}{D}{P^{\dagger}_{\pi}}{U_{1}^{\dagger}}$, we conclude
$\text{Tr}({\texttt{L}_{2}}{U}\rho({t_{0}}){U^{\dagger}})$ $=\text{Tr}({\texttt{L}^{\prime}_{2}}{P_{\pi}}{D}{P_{\pi}^{\dagger}})$
$=\langle{l_{0}}|{P_{\pi}}{D}{P_{\pi}^{\dagger}}|{j_{0}}\rangle$
$=0\penalty 10000\ ,$ (10)
where we used the fact that the permutation matrix simply rearranges the diagonal entries of the matrix $D$. The result in Eq. (III) holds for any permutation matrix $P_{\pi}$} (or even for a composition of permutation matrices). This approach is useful to eliminate all d(d1)d(d-1) eigenmatrices with complex eigenvalues of the Liouvillean that can be triangularized under the unitary mapping ${\texttt{L}_{s}}={U_{1}}{\texttt{L}^{\prime}_{s}}{U_{1}^{\dagger}}$. In other words, such framework is not restricted to the eigenmatrix $\texttt{L}_{2}$}, and is therefore robust enough to eliminate other excited modes that could hinder the exponential acceleration in the relaxation mechanism.
Below, we present the proof that supports our claims. For our purposes, from now on we recast the eigenstates of the Hamiltonian to be written as $|{\psi_{s}}\rangle={U_{1}}|{s}\rangle$, for all $s=\{1,\ldots,d\}$, where $\{|{s}\rangle\}_{s=1,\ldots,d}$}. In this case, the jump operators in Eq. (3) can be conveniently written as ${L_{nm}^{(1)}}={\sqrt{\xi_{nm}}}\,{U_{1}}|{n}\rangle\langle{m}|{U_{1}^{\dagger}}$ and ${L_{nm}^{(2)}}=\sqrt{\chi_{nm}}\,{U_{1}}|{m}\rangle\langle{n}|{U_{1}^{\dagger}}$. The Liouvillian given in Eq. (1) is then vectorized by the Choi-Jamiołkowski isomorphism Jamiołkowski (1972); Choi (1975), which yields
${\mathcal{L}_{\text{vec}}}=({U^{*}_{1}}\otimes{U_{1}})\,\Xi\,({U^{\top}_{1}}\otimes{U^{\dagger}_{1}})\penalty 10000\ ,$ (11)
with
$\Xi:={A^{\dagger}}\otimes{I}+{I}\otimes{A}+{B}\penalty 10000\ ,$ (12)
where $A$ and $B$ are the auxiliary matrices given by
$A=-\frac{1}{2}\,{\sum_{\begin{subarray}{c}n,m=1\\ m<n\end{subarray}}^{d}}\left({\xi_{nm}}|{m}\rangle\langle{m}|+{\chi_{nm}}|{n}\rangle\langle{n}|\right)-i\varepsilon\penalty 10000\ ,$ (13)
and
$B={\sum_{\begin{subarray}{c}n,m=1\\ m<n\end{subarray}}^{d}}\left({\chi_{nm}}|{m,m}\rangle\langle{n,n}|+{\xi_{nm}}|{n,n}\rangle\langle{m,m}|\right)\penalty 10000\ ,$ (14)
with $|m,n\rangle=|{m}\rangle\otimes|{n}\rangle$. Here, ${A^{\dagger}}\otimes{I}+{I}\otimes{A}$ is a diagonal matrix, while $B$ is a non-diagonal matrix that has d(d1)d(d-1) rows/columns with zero elements, since $m<n$.
The operator $\Xi$ in Eq. (12) can be recast into a block-diagonal form, $\Xi={\Xi_{\text{diag}}}\oplus\,{\Xi_{\text{off}}}$, where $\Xi_{\text{diag}}$} is a diagonal matrix with d(d1)d(d-1) rows/columns, while $\Xi_{\text{off}}$} is an off-diagonal matrix with $d$ rows/columns. In addition, it has a spectral decomposition given by $\Xi={\sum_{k}}\,{\lambda_{k}}|{\texttt{R}^{\prime}_{k}}\rangle\rangle\langle\langle{\texttt{L}^{\prime}_{k}}|$, where left $|{\texttt{L}^{\prime}_{k}}\rangle\rangle$ and right $|{\texttt{R}^{\prime}_{k}}\rangle\rangle$ eigenvectors are biorthogonal each other as $\langle\langle{\texttt{L}^{\prime}_{j}}|{\texttt{R}^{\prime}_{k}}\rangle\rangle={\delta_{jk}}$, for all $j,k=\{1,\ldots,{d^{2}}\}$. In this case, it follows that the spectral decomposition of the Liouvillean in Eq. (B5) is written as ${\mathcal{L}_{\text{vec}}}={\sum_{k}}\,{\lambda_{k}}|{\texttt{R}_{k}}\rangle\rangle\langle\langle{\texttt{L}_{k}}|$, where we define $|{X_{k}}\rangle\rangle:=({U^{*}_{1}}\otimes{U_{1}})|{X^{\prime}_{k}}\rangle\rangle$, for all $X\in\{\texttt{L},\texttt{R}\}$. We emphasize that $|{X_{k}}\rangle\rangle$ and $|{X^{\prime}_{k}}\rangle\rangle$ stand for the vector forms of eigenmatrices $X^{\prime}_{k}$ and ${X_{k}}={U_{1}}{X^{\prime}_{k}}{U_{1}^{\dagger}}$, respectively. In addition, we have that vectors $|{\texttt{L}_{j}}\rangle\rangle$ and $|{\texttt{R}_{j}}\rangle\rangle$ satisfies the orthogonality constraint $\langle\langle{\texttt{L}_{j}}|{\texttt{R}_{l}}\rangle\rangle=\langle\langle{\texttt{L}^{\prime}_{j}}|({U_{1}^{\top}}\otimes{U^{\dagger}_{1}})({U^{*}_{1}}\otimes{U_{1}})|{\texttt{R}^{\prime}_{l}}\rangle\rangle=\langle\langle{\texttt{L}^{\prime}_{j}}|{\texttt{R}^{\prime}_{l}}\rangle\rangle={\delta_{jl}}$.
We observe that the operators $\mathcal{L}_{\text{vec}}$} and $\Xi$ share the same spectrum of $d^{2}$} eigenvalues, of which $d$ are real and come from the subblock $\Xi_{\text{off}}$}, and the remaining d(d1)d(d-1) eigenvalues are complex and related to $\Xi_{\text{diag}}$}. On the one hand, we find that the subblock $\Xi_{\text{off}}$} contributes with $d$ eigenvectors $|{\texttt{L}^{\prime}_{s}}\rangle\rangle$ and $|{\texttt{R}^{\prime}_{s}}\rangle\rangle$, whose matrix forms are Hermitian, fully diagonal eigenmatrices $\texttt{L}^{\prime}_{s}$} and $\texttt{R}^{\prime}_{s}$} respective to the basis $\{|{k}\rangle\}_{k=1,\ldots,d}$. This naturally includes the eigenoperator $\texttt{R}^{\prime}_{1}$} related to the steady state ${\texttt{R}_{1}}={U_{1}}{\texttt{R}^{\prime}_{1}}{U_{1}^{\dagger}}$ of the dynamics, and also the vector $|{\texttt{L}^{\prime}_{1}}\rangle\rangle$ constrained to the eigenmatrix $\texttt{L}_{1}$} that plays the role of the identity. On the other hand, the slowest decaying modes of the nonunitary dynamics are related to the subspace spanned by d(d1)d(d-1) eigenvectors of $\Xi_{\text{diag}}$}. Noteworthy, the respective set of left $|{\texttt{L}^{\prime}_{s}}\rangle\rangle$ and right $|{\texttt{R}^{\prime}_{s}}\rangle\rangle$ eigenvectors gives rise to upper/lower triangular matrices $\texttt{L}^{\prime}_{k}$ and $\texttt{R}^{\prime}_{s}$}, respectively, which have all entries equal to zero except for one off-diagonal element. In detail, one finds that $\langle{j}|{\texttt{L}^{\prime}_{s}}|{l}\rangle={\delta_{j,{j_{0}}}}{\delta_{l,{l_{0}}}}$, and $\langle{j}|{\texttt{R}^{\prime}_{s}}|{l}\rangle={\delta_{j,{j_{0}}}}{\delta_{l,{l_{0}}}}$, for a given pair ${j_{0}},{l_{0}}\in\{1,\ldots,d\}$, with ${j_{0}}\neq{l_{0}}$. In particular, we note that this set comprises the eigenoperator $\texttt{L}^{\prime}_{2}$} connected to the lowest decaying mode ${\texttt{L}_{2}}={U_{1}}{\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}$ of the Davies map. It turns out that ${\texttt{L}^{\prime}_{2}}=|{j_{0}}\rangle\langle{l_{0}}|$ is a non-Hermitian, upper/lower triangular matrix.
IV Quantum Mpemba effect
In general, suppressing slowest decaying modes of the dynamics ensures an exponential speed up in the relaxation of the open quantum system. However, this is not a sufficient criterion for the occurrence of a genuine quantum Mpemba effect (QME). The general idea is to identify situations in which a quantum system initially further from equilibrium approaches it more rapidly than an initial state closer to it. To this end, an information-theoretic measure $\mathcal{D}(x,y)$ is introduced to monitor the nonunitary dynamics of the quantum system, thus investigating the distinguishability of the instantaneous state $\rho(t)$ and the steady state $\texttt{R}_{1}$. The genuine QME is evidenced by the crossover between relaxation curves related to initial and transformed states. Useful figures of merit include the Hilbert-Schmidt distance Carollo et al. (2021), the quantum relative entropy Moroder et al. (2024), and the trace distance Liu and Wang , to name a few.
Here, we address the quantum Mpemba effect for open quantum systems whose effective dynamics is governed by Davies maps. Our main result is that the framework described in Sec. III naturally contributes to the occurrence of QME, since (i) it eliminates slower decay modes to accelerate relaxation towards equilibrium; (ii) it ensures that the dressed state ${\rho^{\prime}}({t_{0}})={U}\rho({t_{0}}){U^{\dagger}}$ is as far away from equilibrium as possible compared to the initial state ${\rho}({t_{0}})$, where $U={U_{1}}{P_{\pi}}{\Lambda^{\dagger}}$. This last requirement means that, for a given distinguishability measure $\mathcal{D}(x,y)$, the distance from ${\rho^{\prime}}({t_{0}})$ to the equilibrium is greater than the distance from $\rho({t_{0}})$ to the steady state $\texttt{R}_{1}$}, that is,
$\mathcal{D}({\rho^{\prime}}({t_{0}}),{\texttt{R}_{1}})>\mathcal{D}(\rho({t_{0}}),{\texttt{R}_{1}})\penalty 10000\ .$ (15)
We note that ${\texttt{R}_{1}}={U_{1}}\Sigma{U_{1}^{\dagger}}$ defines the steady state of the system, with $\Sigma=\exp(-\varepsilon/{k_{B}}T)/Z$, and $Z=\text{Tr}(\exp(-\varepsilon/{k_{B}}T))$ is the partition function. Once conditions (i) and (ii) are fully satisfied, a genuine QME occurs if, at some later time ${t_{\text{QME}}}>{t_{0}}$, the following reverse bound is found
$\mathcal{D}({\rho^{\prime}}(t),{\texttt{R}_{1}})<\mathcal{D}(\rho(t),{\texttt{R}_{1}})\penalty 10000\ ,$ (16)
for all $t>{t_{\text{QME}}}$, where ${\rho^{\prime}}(t)={e^{t\mathcal{L}}}[{\rho^{\prime}}({t_{0}})]$ and $\rho(t)={e^{t\mathcal{L}}}[{\rho}({t_{0}})]$ define the instantaneous states obtained from ${\rho^{\prime}}({t_{0}})$ and $\rho({t_{0}})$, respectively, both evolving under the generator in Eq. (1).
To be valid, the bound in Eq. (16) requires that the two curves of $\mathcal{D}({\rho^{\prime}}(t),{\texttt{R}_{1}})$ and $\mathcal{D}(\rho(t),{\texttt{R}_{1}})$ intersect at time $t_{\text{QME}}$}. This crossover remains as a consequence of points (i) and (ii) to be satisfied for the proposed unitary transformation. We already proven point (i), that is, the unitary $U$ contributes to speed up relaxation. In the following, we verify point (ii) by showing that the bound in Eq. (15) is valid for three paradigmatic distinguishability measures, namely, Hilbert-Schmidt distance, quantum relative entropy, and trace distance. In detail, we prove that there always exists a permutation matrix $P_{\pi}$} that maximizes the distance $\mathcal{D}({\rho^{\prime}}({t_{0}}),{\texttt{R}_{1}})$ from the dressed state to the equilibrium, compared to $\mathcal{D}({\rho}({t_{0}}),{\texttt{R}_{1}})$. Some technical details of the proofs can be found in Appendix B.
IV.1 Hilbert-Schmidt distance
Let $\rho\in\mathcal{S}$ be two density matrices defined on the convex space of quantum states $\mathcal{S}=\{\rho\in\mathcal{H}\penalty 10000\ |\penalty 10000\ \rho^{\dagger}=\rho,\penalty 10000\ \rho\geq 0,\penalty 10000\ \text{Tr}(\rho)=1\}$, where $\mathcal{H}$ is a $d$-dimensional Hilbert space. The Hilbert-Schmidt distance (HSD) between these two states is defined as
${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)=\sqrt{\text{Tr}(|\rho-\varrho|)}\penalty 10000\ ,$ (17)
with $|{X}|:=\sqrt{{X^{\dagger}}X}$. We note that HSD is (i) non-negative, i.e., ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)\geq 0$, with ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)={\mathcal{D}_{\text{HSD}}}(\varrho,\rho)$; (iii) isometric invariant, ${\mathcal{D}_{\text{HSD}}}(V\rho{V^{\dagger}},V\varrho{V^{\dagger}})={\mathcal{D}_{\text{HSD}}}(\rho,\varrho)$, where ${V^{\dagger}}={V^{-1}}$ is a unitary operator. It is worth to note that HSD satisfy the triangle inequality, ${\mathcal{D}_{\text{HSD}}}(\rho,\varrho)\leq{\mathcal{D}_{\text{HSD}}}(\rho,\varpi)+{\mathcal{D}_{\text{HSD}}}(\varpi,\varrho)$, for all $\rho,\varpi,\varrho\in\mathcal{S}$. However, HSD it is not always contractive under completely positive and trace-preserving maps Ozawa (2000).
The bound in Eq. (15), applied to the Hilbert-Schmidt distance, becomes ${\mathcal{D}_{\text{HSD}}}({\rho^{\prime}}({t_{0}}),{\texttt{R}_{1}})>{\mathcal{D}_{\text{HSD}}}({\rho}({t_{0}}),{\texttt{R}_{1}})$. This inequality, when satisfied, must imply the following upper bound
$\text{Tr}({\rho^{\prime}}({t_{0}}){\texttt{R}_{1}})<\text{Tr}({\rho}({t_{0}}){\texttt{R}_{1}})\penalty 10000\ .$ (18)
In Appendix B, we show that Eq. (18) can be written as
$\text{Tr}({P_{\pi}}{D}{P^{\dagger}_{\pi}}{\Sigma})<{\sum_{l}}\,{\epsilon_{l}}\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)\penalty 10000\ ,$ (19)
where $A_{l}$ is a permutation matrix, $0\leq{\epsilon_{l}}\leq 1$, with ${\sum_{l}}\,{\epsilon_{l}}=1$, while $D$ and $\Sigma$ are diagonal matrices containing the spectrum of $\rho({t_{0}})$ and $\texttt{R}_{1}$}, respectively. The right-hand side of Eq. (19) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{c_{l}}$ of the nonnegative elements ${c_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)$. We note that ${c_{\text{min}}}\leq{\sum_{l}}\,{\epsilon_{l}}{c_{l}}\leq{c_{\text{max}}}$, where ${c_{\text{min}}}=\text{min}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$ and ${c_{\text{max}}}=\text{max}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$. This means that there exist an optimal permutation matrix $P_{\pi,\text{opt}}$} such that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)$ is closest to the minimum value $c_{\text{min}}$, i.e., $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$. To see this point, we note that
$\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)={\sum_{l=1}^{d}}\,{\alpha_{l}}{\lambda_{\pi(l)}}\penalty 10000\ ,$ (20)
where $\pi(\bullet)$ maps index $l$ to $\pi(l)$, with $l,\pi(l)\in\{1,2,\ldots,d\}$. Therefore, to ensure that ${\sum_{l=1}^{d}}\,{\alpha_{l}}{\lambda_{\pi(l)}}\leq{c_{\text{min}}}$, it suffices to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\leq{\lambda_{\pi(2)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{n}}$ are listed in descending order, or vice-versa Hardy et al. (1988). This proves the validity of the inequality in Eq. (18), which means that there exists a permutation $P_{\pi}$ that maximizes HSD evaluated for states ${\rho^{\prime}}({t_{0}})$ and $\texttt{R}_{1}$}.
IV.2 Quantum relative entropy
The quantum relative entropy (QRE) respective to states $\rho,\varrho\in\mathcal{S}$ is defined as
${\mathcal{D}_{\text{QRE}}}(\rho,\varrho):=-S(\rho)-\text{Tr}(\rho\ln\varrho)\penalty 10000\ ,$ (21)
where $S(\rho)=-\text{Tr}(\rho\ln\rho)$ is the von Neumann entropy Umegaki (1962). To be clear, QRE is (i) non-negative, i.e., ${\mathcal{D}_{\text{QRE}}}(\rho,\varrho)\geq 0$, with ${\mathcal{D}_{\text{QRE}}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) isometric invariant, ${\mathcal{D}_{\text{QRE}}}(V\rho{V^{\dagger}},V\varrho{V^{\dagger}})={\mathcal{D}_{\text{QRE}}}(\rho,\varrho)$, where ${V^{\dagger}}={V^{-1}}$ is a unitary operator; (iii) monotonically decreasing under completely positive and trace preserving (CPTP) maps, i.e., ${\mathcal{D}_{\text{QRE}}}\left(\mathcal{E}(\rho),\mathcal{E}(\varrho)\right)\leq{\mathcal{D}_{\text{QRE}}}(\rho,\varrho)$, with $\mathcal{E}(\bullet)$ being a given CPTP operation Vedral (2002); Schumacher and Westmoreland .
The bound in Eq. (16), applied to QRE, becomes ${\mathcal{D}_{\text{QRE}}}({\rho^{\prime}}({t_{0}}),{\texttt{R}_{1}})>{\mathcal{D}_{\text{QRE}}}({\rho}({t_{0}}),{\texttt{R}_{1}})$. This inequality, when satisfied, readily implies the lower bound
$\text{Tr}({\rho^{\prime}}({t_{0}})H)>\text{Tr}({\rho}({t_{0}})H)\penalty 10000\ ,$ (22)
where we used the fact that the von Neumann entropy is unitarily invariant, i.e., $S({\rho^{\prime}}({t_{0}}))=S(U\rho({t_{0}}){U^{\dagger}})=S({\rho}({t_{0}}))$. We also used that $\text{Tr}(X\ln{\texttt{R}_{1}})=-\text{Tr}(XH)-\ln{Z}$, which holds for $X=\rho({t_{0}})$, and $X={\rho^{\prime}}({t_{0}})$. Next, we recall that ${\rho^{\prime}}({t_{0}})={U_{1}}{P_{\pi}}{D}{P_{\pi}^{\dagger}}{U_{1}^{\dagger}}$, and $H={U_{1}}\varepsilon{U_{1}^{\dagger}}$, so that
$\text{Tr}({\rho^{\prime}}({t_{0}})H)=\text{Tr}({P_{\pi}}D{P_{\pi}^{\dagger}}\varepsilon)\penalty 10000\ .$ (23)
Based on the results from Appendix B, it can be verify that
$\text{Tr}({\rho}({t_{0}})H)={\sum_{l}}\,{\epsilon_{l}}\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\varepsilon)\penalty 10000\ ,$ (24)
where $A_{l}$ is a permutation matrix, with $0\leq{\epsilon_{l}}\leq 1$, with ${\sum_{l}}\,{\epsilon_{l}}=1$. Hence, by combining Eqs. (22), (23), and (24), it yields
$\text{Tr}({P_{\pi}}{D}{P_{\pi}^{\dagger}}\varepsilon)>{\sum_{l}}\,{\epsilon_{l}}\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\varepsilon)\penalty 10000\ .$ (25)
The right-hand side of Eq. (25) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{g_{l}}$ of the nonnegative elements ${g_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\varepsilon)$. To validate the inequality in Eq. (25), we apply the same reasoning as in Sec. IV.2. The ideia is that there exist an optimal permutation matrix $P_{\pi,\text{opt}}$} that satisfies the lower bound $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\geq{g_{\text{max}}}\geq{\sum_{l}}\,{\epsilon_{l}}{g_{l}}$, where ${g_{\text{max}}}=\text{max}\{{g_{1}},{g_{2}},\ldots,{g_{d}}\}$. We note that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)={\sum_{l}}\,{\alpha_{l}}{\lambda_{\pi(l)}}$, where $\pi(\bullet)$ maps index $l$ to $\pi(l)$, with $l,\pi(l)\in\{1,2,\ldots,d\}$. Therefore, to ensure that ${\sum_{l}}\,{\alpha_{l}}{\lambda_{\pi(l)}}\geq{g_{\text{max}}}\geq{\sum_{l}}\,{\epsilon_{l}}{g_{l}}$, it is sufficient to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\geq{\lambda_{\pi(2)}}\geq\ldots\geq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{n}}$ are listed in ascending order, or vice-versa Hardy et al. (1988). This proves the inequality in Eq. (22), which means that there exists $P_{\pi}$ that maximizes QRE evaluated for ${\rho^{\prime}}({t_{0}})$ and $\texttt{R}_{1}$}.
IV.3 Trace distance
The trace distance (TD) for two quantum states $\rho,\varrho\in\mathcal{S}$ is defined as
${\mathcal{D}_{\text{TD}}}(\rho,\varrho)=\frac{1}{2}\text{Tr}(|\rho-\varrho|)\penalty 10000\ ,$ (26)
with $|A|=\sqrt{{A^{\dagger}}A}$. Overall, TD is (i) non-negative, i.e., ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)\geq 0$, with ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)={\mathcal{D}_{\text{TD}}}(\varrho,\rho)$; (iii) unitarily invariant, ${\mathcal{D}_{\text{TD}}}(V\rho{V^{\dagger}},V\varrho{V^{\dagger}})={\mathcal{D}_{\text{TD}}}(\rho,\varrho)$, where ${V^{\dagger}}={V^{-1}}$ is a unitary operator; (iv) monotonically decreasing under completely positive and trace preserving (CPTP) maps, i.e., ${\mathcal{D}_{\text{TD}}}\left(\mathcal{E}(\rho),\mathcal{E}(\varrho)\right)\leq{\mathcal{D}_{\text{TD}}}(\rho,\varrho)$, with $\mathcal{E}(\bullet)$ being a given CPTP operation Nielsen and Chuang (2010).
Figure 2: (Color online) The quantum Mpemba effect in the dynamics of the transverse field Ising (TFI) model with open boundary conditions [see details in Sec. V], where $N=5$ spins, $h=J/2$, $\gamma=1$, and $T=0.1$. To investigate the relaxation of the instantaneous state $\rho(t)$ towards the steady state $\texttt{R}_{1}$}, we consider the following figures of merit: (a) Hilbert-Schmidt distance, ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})$; (b) quantum relative entropy, ${\mathcal{D}_{\text{QRE}}}(\rho(t),{\texttt{R}_{1}})$; (c) trace distance, ${\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})$. In each panel, we plot each of these distinguishability measures as a function of the dimensionless parameter $Jt$. The blue solid lines refer to the initial state $\rho({t_{0}})$ of the system, while the dressed state ${\rho^{\prime}}({t_{0}})$ is represented by red dashed lines.
To investigate QME, we first note that the trace distance for the initial state $\rho({t_{0}})$ and the steady state $\texttt{R}_{1}$ satisfies the lower bound Markham et al. (2008)
${\mathcal{D}_{\text{TD}}}({u^{\uparrow}},{v^{\downarrow}})\geq{\mathcal{D}_{\text{TD}}}(\rho({t_{0}}),{\texttt{R}_{1}})\penalty 10000\ ,$ (27)
where ${u^{\uparrow}}=\text{diag}({\lambda_{1}},\ldots,{\lambda_{d}})$ defines a vector formed by the eigenvalues of $\rho({t_{0}})$ listed in ascending order, while ${v^{\downarrow}}=\text{diag}({q_{1}},\ldots,{q_{d}})$ is a vector that contains the eigenvalues of $\texttt{R}_{1}$ listed in descending order, with ${q_{l}}=e^{-{\varepsilon_{l}}/{k_{B}}T}/Z$ being the $l$-th eigenvalue of the steady state. Next, note that the trace distance for the initial dressed state ${\rho^{\prime}}({t_{0}})$ and the steady state $\texttt{R}_{1}$ is written as
${\mathcal{D}_{\text{TD}}}({\rho^{\prime}}({t_{0}}),{\texttt{R}_{1}})$ $=\frac{1}{2}\text{Tr}\left[\sqrt{({U_{1}}{P_{\pi}}{D}{P_{\pi}^{\dagger}}{U_{1}^{\dagger}}-{U_{1}}\Sigma{U_{1}^{\dagger}})^{2}}\,\right]$
$=\frac{1}{2}\text{Tr}\left[\sqrt{({P_{\pi}}{D}{P_{\pi}^{\dagger}}-\Sigma)^{2}}\,\right]$
$={\mathcal{D}_{\text{TD}}}({P_{\pi}}{D}{P_{\pi}^{\dagger}},\Sigma)\penalty 10000\ ,$ (28)
where we used the fact the TD is unitarily invariant. Here, $\Sigma=\text{diag}({q_{1}},\ldots,{q_{d}})$, while ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},\ldots{\lambda_{\pi(d)}})$ is the matrix obtained by rearranging the entries of the diagonal matrix $D$ that contains the eigenvalues of the initial state. It is worth noting that always exist an appropriate permutation matrix $P_{\pi}$} that rearranges the eigenvalues of the initial state such that ${\mathcal{D}_{\text{TD}}}({P_{\pi}}{D}{P_{\pi}^{\dagger}},\Sigma)={\mathcal{D}_{\text{TD}}}({u^{\uparrow}},{v^{\downarrow}})$. In this case, combining Eqs. (27) and (IV.3), it yields
${\mathcal{D}_{\text{TD}}}({\rho^{\prime}}({t_{0}}),{\texttt{R}_{1}})\geq{\mathcal{D}_{\text{Tr}}}(\rho({t_{0}}),{\texttt{R}_{1}})\penalty 10000\ .$ (29)
Hence, satisfied the inequality in Eq. (29), that is, if the trace distance between dressed initial state and the steady state is maximized for a given permutation matrix, a genuine quantum Mpemba effect is expected to be observed.
V Examples
In the following, we illustrate our findings for QME by investigating the nonunitary dynamics of two paradigmatic spin models with open boundary conditions, namely, the transverse field Ising (TFI) model,
${H_{\text{TFI}}}=-{J}\,{\sum_{j=1}^{N-1}}{\sigma_{j}^{z}}{\sigma_{j+1}^{z}}+{h}\,{\sum_{j=1}^{N}}{\sigma_{j}^{x}}\penalty 10000\ ,$ (30)
and the XXZ model,
${H_{\text{XXZ}}}={\sum_{j=1}^{N-1}}[J({\sigma_{j}^{x}}{\sigma_{j+1}^{x}}+{\sigma_{j}^{y}}{\sigma_{j+1}^{y}})+\Delta{\sigma_{j}^{z}}{\sigma_{j+1}^{z}}]\penalty 10000\ .$ (31)
Here, $\sigma_{j}^{x,y,z}$} denotes the Pauli matrices acting on the $j$-th site of the chains, $N$ is the total number of spins, $J$ is the coupling between neighboring spins, $h$ corresponds to the strength of the transverse magnetic field, and $\Delta$ is the anisotropy parameter. These spin chains are weakly coupled to a bosonic reservoir at temperature $T$. The nonunitary dynamics of each system is governed by the Davies map discussed in Sec. II.
The initial state is defined as $\rho({t_{0}})={U_{1}}|\phi\rangle\langle{\phi}|{U_{1}^{\dagger}}$ at time ${t_{0}}=0$, where $U_{1}$ is the unitary matrix formed by the eigenstates of the Hamiltonian, and $|\phi\rangle=({1}/{\sqrt{d}}){\sum_{l=1}^{d}}\,|{j}\rangle$, while $\{|{j}\rangle\}_{j=1,\ldots,d}$ is the eigenbasis of the collective spin operator ${S_{z}}=(1/2){\sum_{j=1}^{d}}{\sigma_{j}^{z}}$, with $d={2^{N}}$. The permutation matrix ${P_{\pi}}={\sigma_{x}^{\otimes{N}}}$ is chosen to maximize the distinguishability between states ${\rho^{\prime}}({t_{0}})$ and ${\texttt{R}_{1}}={U_{1}}\Sigma{U_{1}^{\dagger}}$, with $\Sigma=\exp(-\varepsilon/{k_{B}}T)/Z$, and $Z=\text{Tr}(\exp(-\varepsilon/{k_{B}}T))$, where $\varepsilon=\text{diag}({\varepsilon_{1}},\ldots,{\varepsilon_{d}})$ encodes the energies of each Hamiltonian, with ${\varepsilon_{1}}>\ldots>{\varepsilon_{d}}$.
Figure 3: (Color online) The quantum Mpemba effect in the dynamics of the XXZ model with open boundary conditions [see details in Sec. V], where $N=5$ spins, $\Delta=J/2$, $\gamma=1$, and $T=0.1$. To investigate the relaxation of the instantaneous state $\rho(t)$ towards the steady state $\texttt{R}_{1}$}, we consider the following figures of merit: (a) Hilbert-Schmidt distance, ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})$; (b) quantum relative entropy, ${\mathcal{D}_{\text{QRE}}}(\rho(t),{\texttt{R}_{1}})$; (c) trace distance, ${\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})$. In each panel, we plot each of these distinguishability measures as a function of the dimensionless parameter $Jt$. The blue solid lines refer to the initial state $\rho({t_{0}})$ of the system, while the dressed state ${\rho^{\prime}}({t_{0}})$ is represented by red dashed lines.
Figures 2 and 3 present the numerical simulations for TFI and XXZ models, respectively. We set $N=5$, $h=J/2$, $\Delta=J/2$, $\gamma=1$, $T=0.1$, and ${k_{B}}=1$. In these figures, we show the plots for Hilbert-Schmidt distance [see Figs. 2(a) and 3(a)], quantum relative entropy [see Figs. 2(b) and 3(b)], and trace distance [see Figs. 2(c) and 3(c)], as functions of the dimensionless parameter $Jt$. The blue solid lines refer to ${\mathcal{D}_{\text{HSD},\text{QRE},\text{TD}}}(\rho(t),{\texttt{R}_{1}})$, with $\rho(t)={e^{t\mathcal{L}}}[{\rho}({t_{0}})]$, while ${\mathcal{D}_{\text{HSD},\text{QRE},\text{TD}}}({\rho^{\prime}}(t),{\texttt{R}_{1}})$ is represented by red dashed lines, with ${\rho^{\prime}}(t)={e^{t\mathcal{L}}}[{\rho^{\prime}}({t_{0}})]$.
The insets in Figs. 2 and 3 show the crossovers that clearly indicate the occurrence of the genuine Mpemba effect for each spin model. For the TFI model, we find the crossover to occur at (i) $J{t_{\text{QME}}}\approx 1.47$ for HSD [see Fig. 2(a)], (ii) $J{t_{\text{QME}}}\approx 6.16$ for QRE [see Fig. 2(b)], and (iii) $J{t_{\text{QME}}}\approx 1.57$ for TD [see Fig. 2(c)]. For the XXZ model, we find the crossover to occur at (i) $J{t_{\text{QME}}}\approx 1.23$ for HSD [see Fig. 3(a)], (ii) $J{t_{\text{QME}}}\approx 7.45$ for QRE [see Fig. 3(b), and (iii) $J{t_{\text{TD}}}\approx 1.85$ [see Fig. 3(c)]. The plots for the Hilbert-Schmidt distance and trace distance exhibit similar behavior when compared to each other, regardless of the spin model, except at the beginning of the nonunitary evolution. Indeed, the insets show that the HSD in Figs. 2(a) and 3(a) exhibit a more pronounced decay than the TD in Figs. 2(c) and 3(c), at earlier times of the dynamics, for both the TFI and XXZ models.
VI Accelerated relaxation and quantum coherences
We proved that the exponential speed up on the dynamics of a given Davies map depends on the fact that the transformed state ${\rho^{\prime}}({t_{0}})=U\rho({t_{0}}){U^{\dagger}}$ is incoherent with respect to the eigenbasis $\{|{\psi_{l}}\rangle\}_{l=1,\ldots,d}$} of the Hamiltonian $H$. This suggests that, for a probe state $\rho({t_{0}})$ that is already incoherent in this preferred basis, one can naturally expect faster relaxation to equilibrium. This property, however, can also be observed for a probe state that exhibit coherence in such a basis. Here we argue that there are initial states which spontaneously lead to exponentially fast relaxation to the equilibrium, but they need not be incoherent in the eigenbasis of the Hamiltonian of the open quantum system.
Let $\rho({t_{0}})$ an initial state of an open quantum system described by a Davies map. This state may or may not be incoherent in the energy eigenbasis. We recall that ${\texttt{L}_{2}}={U_{1}}{\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}$, where ${\texttt{L}^{\prime}_{2}}=|{k_{0}}\rangle\langle{l_{0}}|$ is a non-Hermitian matrix that has a single nonzero element related to the pair $({k_{0}},{l_{0}})\in\{1,2,\ldots,d\}$, with ${k_{0}}\neq{l_{0}}$. Then, the overlap between $\texttt{L}_{2}$} and $\rho({t_{0}})$ becomes
$\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))$ $=\text{Tr}({\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}\rho({t_{0}}){U_{1}})$
$=\langle{l_{0}}|{U_{1}^{\dagger}}\rho({t_{0}}){U_{1}}|{k_{0}}\rangle$
$=\langle{\psi_{l_{0}}}|\rho({t_{0}})|{\psi_{k_{0}}}\rangle\penalty 10000\ ,$ (32)
where we used the fact that $|{\psi_{s}}\rangle={U_{1}}|{s}\rangle$ is the $s$-th eigenstate of $H$.
It is worth noting that Eq. (VI) allows for three possible scenarios. First, a vanishing overlap is obtained whenever $\rho({t_{0}})$ is an incoherent state in the energy eigenbasis. This means that the relaxation to equilibrium is expected to be accelerated. Second, for an initial state with nonzero coherences in this preferred basis, particularly with respect to the states $|{l_{0}}\rangle$ and $|{k_{0}}\rangle$, we obtain $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))=\langle{\psi_{l_{0}}}|\rho({t_{0}})|{\psi_{k_{0}}}\rangle\neq 0$. In this case, faster relaxation can be achieved by using the dressed state ${\rho^{\prime}}({t_{0}})=U\rho({t_{0}}){U^{\dagger}}$ as an input to the nonunitary evolution. Finally, there may exist a probe state $\rho({t_{0}})$ that exhibit nonzero coherences for all eigenstates of the Hamiltonian, except for the vector pair $|{l_{0}}\rangle$ and $|{k_{0}}\rangle$, i.e., $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))=\langle{\psi_{l_{0}}}|\rho({t_{0}})|{\psi_{k_{0}}}\rangle=0$. To see this, it is sufficient that the matrix ${U_{1}^{\dagger}}\rho({t_{0}}){U_{1}}$ has a vanishing off-diagonal element whose position coincides with that one of the single nonzero off-diagonal element of the eigenmatrix $\texttt{L}^{\prime}_{2}$.
To illustrate these findings, we consider a system of two-qubits with Hamiltonian $H={U_{1}}\varepsilon{U_{1}^{\dagger}}$, with $\varepsilon=\text{diag}({\varepsilon_{1}},\ldots,{\varepsilon_{4}})$, where the energies are listed in descending order, ${\varepsilon_{j}}>{\varepsilon_{j+1}}$ for $j=\{1,2,3\}$. Here, ${U_{1}}=|{0,+}\rangle\langle{0,0}|+|{0,-}\rangle\langle{0,1}|+|{1,0}\rangle\langle{1,0}|+|{1,1}\rangle\langle{1,1}|$ is the unitary that brings the Hamiltonian of the system into its diagonal form, with $|{\pm}\rangle=(1/\sqrt{2})(|{0}\rangle\pm|{1}\rangle)$. Let ${\texttt{L}^{\prime}_{2}}=|{0,1}\rangle\langle{0,0}|$ be the eigenoperator related to the slowest decaying mode of the Liouvillian. On the one hand, for the initial state $\rho({t_{0}})=|{0,0}\rangle\langle{0,0}|$, we have that ${U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}=|{0,+}\rangle\langle{0,+}|$, and one obtains the nonzero off-diagonal element $\langle{0,0}|{U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}|{0,1}\rangle=1/2$. In this case, the dynamics of the quantum system can be exponentially accelerated by applying a suitable unitary matrix on the probe state such that it becomes diagonal respective to the eigenbasis of the Hamiltonian. In this case, to speed up the relaxation process, we follow our protocol and map $\rho({t_{0}})$ into the dressed state $\rho({t_{0}})$. On the other hand, when choosing the probe state $\rho({t_{0}})=|{1,+}\rangle\langle{1,+}|$, one obtains ${U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}=|{1,+}\rangle\langle{1,+}|$, which implies a zero valued off-diagonal element $\langle{1,0}|{U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}|{1,1}\rangle=0$. Hence, it follows that the dynamics of the quantum system is already exponentially accelerated, even thought $\rho({t_{0}})$ is not incoherent with respect to the energy eigenbasis.
VII Conclusions
In conclusion, we have presented a framework to generate a genuine quantum Mpemba effect in the relaxation of open quantum systems. The protocol implements unitary operations built from permutation matrices that cause a rearrangement of the spectrum of a given initial state. On the one hand, the slower decay mode of the Liouville operator is suppressed, which triggers an exponentially faster relaxation to the equilibrium. On the other hand, such a protocol maps an input state to an incoherent state in the energy eigenbasis that is as far away as possible from the steady state of the dynamics. Once these conditions are met, we guarantee that a genuine Mpemba effect will occur.
To characterize the relaxation process, we monitor the distinguishability between instantaneous states and the steady state of the open quantum system. To do so, we consider Hilbert-Schmidt distance, quantum relative entropy, and trace distance. We have shown that, for any initial state, there always exists a permutation matrix that maximizes the distance from the dressed state to equilibrium. Compared to the results of Ref. Moroder et al. (2024), our protocol is implementation-friendly since its numerical simulation requires low computational effort. We emphasize that our framework is capable of retrieving previous results discussed in the literature, making appropriate modifications and refining some assumptions. Our results contribute to elucidate the mechanism underlying the quantum Mpemba effect, and pave the way for insightful perspectives for the study of the dynamics of open quantum systems.
Acknowledgements.
This work was supported by the Brazilian ministries MEC and MCTIC, and the Brazilian funding agencies CNPq, and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior–Brasil (CAPES) (Finance Code 001). D. P. P. would like to acknowledge the Fundação de Amparo à Pesquisa e ao Desenvolvimento Científico e Tecnológico do Maranhão (FAPEMA) (Grant No. PQ-C-12651/25).
Appendix
A Exponentially accelerated relaxation for real slowest decaying mode
Here, we show that the results from Sec. III can be properly modified to retrieve those addressed in Ref. Carollo et al. (2021). We consider the dynamics of an open quantum system coupled to a Markovian environment described by the Markovian master equation $d\rho(t)/dt=\mathcal{L}[\rho(t)]$, where $\rho(t)$ is the instantaneous state of the system for all $t\geq 0$, while $\mathcal{L}[\bullet]$ is the Lindblad superoperator as follows Breuer and Petruccione (2007)
$\mathcal{L}[\bullet]=-i[H,\bullet]+{\sum_{j=1}^{N}}\left({L_{j}}\bullet{L^{\dagger}_{j}}-\frac{1}{2}\left\{{L^{\dagger}_{j}}{L_{j}},\bullet\right\}\right)\penalty 10000\ .$ (A1)
[…]
Here, we set the lowest decaying mode $\texttt{L}_{2}$} of the Liouvillian to be Hermitian, i.e., ${\texttt{L}^{\dagger}_{2}}={\texttt{L}_{2}}$, with $\lambda_{2}$ being a real eigenvalue. On the other hand, the Hermitian eigenmatrix admits the Schur decomposition ${\texttt{L}_{2}}={U_{1}}{\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}$, where ${\texttt{L}^{\prime}_{2}}=\text{diag}({\alpha_{1}},\ldots,{\alpha_{d}})$ is a diagonal matrix that takes into account the set of real eigenvalues of $\texttt{L}_{2}$}, and $U_{1}$} is a unitary matrix formed by the respective eigenstates Artin (2017). On the one hand, the pure initial state of the system is written as $\rho({t_{0}})=\Lambda{D}{\Lambda^{\dagger}}$, where $D=\text{diag}(0,\ldots,1,\ldots,0)$ is a diagonal matrix whose elements are the eigenvalues of the probe state, while $\Lambda$ is a unitary matrix formed by the respective eigenstates.
Next, we consider the probe state $\rho({t_{0}})$ is mapped to ${\rho^{\prime}}(t_{0})=U\rho({t_{0}}){U^{\dagger}}$ by means of the unitary operator $U={U_{1}}V(\theta){\Lambda^{\dagger}}$. Here, we have appropriately choose $V(\theta)$ as the unitary matrix as follows
$V(\theta)={e^{i\theta{S_{\pi}}}}{P_{\pi}}=\left(\cos\theta\,{\mathbb{I}_{d}}+i\sin\theta{S_{\pi}}\right){P_{\pi}}\penalty 10000\ ,$ (A2)
with $\theta\in[0,\pi/2]$, while $S_{\pi}$} and $P_{\pi}$} are real-valued permutation matrices. We note that ${S_{\pi}^{\top}}={S_{\pi}}$ is a symmetric matrix, while ${P_{\pi}^{\dagger}}={P_{\pi}^{-1}}$ is a unitary permutation matrix. In this setting, the overlap between the eigenmatrix $\texttt{L}_{2}$} and the transformed state $U\rho({t_{0}}){U^{\dagger}}$ is written as
$\text{Tr}\left({\texttt{L}_{2}}{U}\rho({t_{0}}){U^{\dagger}}\right)={\cos^{2}}\theta\,\text{Tr}({\texttt{L}^{\prime}_{2}}{P_{\pi}}{D}{P_{\pi}^{\dagger}})$
$+{\sin^{2}}\theta\,\text{Tr}({P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}{S_{\pi}}{D}{S_{\pi}^{\dagger}})$
$+\frac{i}{2}\sin(2\theta)\text{Tr}\left({S_{\pi}}[{D},{P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}]\right)\penalty 10000\ .$ (A3)
In the following, we will discuss how the elimination of the slowest decay mode can be achieved.
Let $P_{\pi}$} be an arbitrary permutation matrix. We note that the diagonal matrix $\texttt{L}^{\prime}_{2}$} is always mapped onto another diagonal matrix ${P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}$ under the action of some permutation matrix $P_{\pi}$}. The overall effect is the rearrangement of the entries along the diagonal of the original matrix. Therefore, since the diagonal matrices $D$ and ${P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}$ commutes each other, i.e., $D{P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}={P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}D$, one can readily concludes that
$\text{Tr}({S_{\pi}}[D,{P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}])=0\penalty 10000\ ,$ (A4)
regardless of the permutation matrix chosen. The remaining coefficients $\text{Tr}({\texttt{L}^{\prime}_{2}}{P_{\pi}}{D}{P_{\pi}^{\dagger}})$ and $\text{Tr}({P_{\pi}^{\dagger}}{\texttt{L}^{\prime}_{2}}{P_{\pi}}{S_{\pi}}{D}{S_{\pi}^{\dagger}})$ in the right-hand side of Eq. (A) are not necessarily equal to zero. In fact, they are related to the eigenvalues of the operator $\texttt{L}_{2}$} and, for a given critical value $\theta_{c}$, they both can combine and cancel each other out, thus ensuring the overlap in Eq. (A) is zero. To outline this proof, we follow here a similar technical discussion to that presented in Ref. Carollo et al. (2021). We note that, for a suitable choice for the permutation matrix $P_{\pi}$}, we have that the diagonal matrix $D$ is mapped onto ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}(1,0,\ldots,0)$, which in turn implies that
$\text{Tr}({\texttt{L}^{\prime}_{2}}{P_{\pi}}{D}{P_{\pi}^{\dagger}})={\alpha_{1}}\penalty 10000\ .$ (A5)
To address the overlap $\text{Tr}({P_{\pi}^{\dagger}}{T}{P_{\pi}}{S_{\pi}}{D}{S_{\pi}^{\dagger}})$, two points need to be highlighted. First, since the left eigenmatrix is Hermitian, it can be written as ${\texttt{L}_{2}}={\sum_{k=1}^{d}}\,{\alpha_{k}}|{\varphi_{k}}\rangle\langle{\varphi_{k}}|$, where $\{|{\varphi_{k}}\rangle\}_{k=1,\ldots,d}$} is the set of eigenstates of $\texttt{L}_{2}$} that defines a basis for the Hilbert space, with $\langle{\varphi_{k}}|{\varphi_{l}}\rangle={\delta_{kl}}$, and ${\sum_{k=1}^{d}}\,|{\varphi_{k}}\rangle\langle{\varphi_{k}}|={\mathbb{I}_{d}}$. Secondly, we recall that the eigenmatrices $\texttt{R}_{1}$ and $\texttt{L}_{2}$ are normalized to be biorthogonal each other, that is, $\text{Tr}({\texttt{R}_{1}}{\texttt{L}_{2}})=0$. Hence, we obtain the constraint as follows
${\sum_{k=1}^{d}}\,{\alpha_{k}}\langle{\varphi_{k}}|{\texttt{R}_{1}}|{\varphi_{k}}\rangle=0\penalty 10000\ .$ (A6)
The right eigenmatrix $\texttt{R}_{1}$} is the steady state of the system, i.e., it is a positive-semidefinite matrix that satisfies $\langle{\varphi_{k}}|{\texttt{R}_{1}}|{\varphi_{k}}\rangle\geq 0$ for any eigenstate $|{\varphi_{k}}\rangle$ of the operator $\texttt{L}_{2}$. This means that, to satisfy Eq. (A6), the set of eigenvalues cannot consist exclusively by positive real numbers, nor only of negative real numbers. In fact, there must be at least two nonzero eigenvalues with opposite sign, for example, the two eigenvalues $\alpha_{1}$ and $\alpha_{n}$ such that $\text{sign}({\alpha_{n}})=-\text{sign}({\alpha_{1}})$. We then construct permutation matrices $P_{\pi}$} and $S_{\pi}$} to rearrange the ordering of the diagonal entries of matrices $T$ and $D$, respectively, so that
$\text{Tr}({P_{\pi}^{\dagger}}{T}{P_{\pi}}{S_{\pi}}{D}{S_{\pi}^{\dagger}})={\alpha_{n}}\penalty 10000\ .$ (A7)
Therefore, combining Eq. (A) with Eqs. (A4), (A5), and (A7), we obtain the result as follows
$\text{Tr}\left({\texttt{L}_{2}}{U}\rho({t_{0}}){U^{\dagger}}\right)={\alpha_{1}}{\cos^{2}}\theta+{\alpha_{n}}{\sin^{2}}\theta\penalty 10000\ .$ (A8)
The slowest decaying mode can be neglected if, for a given critical angle $\theta_{c}$, we have that $\text{Tr}\left({\texttt{L}_{2}}{U}\rho({t_{0}}){U^{\dagger}}\right)=0$, which requires that
${\theta_{c}}=\arctan\left(\sqrt{\left|\frac{\alpha_{1}}{\alpha_{n}}\right|}\,\right)\penalty 10000\ .$ (A9)
We note that Eqs. (A8) and (A9) exactly recover the results obtained in Ref. Carollo et al. (2021).
B Proof of Eq. (19)
In this Appendix, we present the proof of Eq. (19) given in the main text of the manuscript. The Hilbert-Schmidt distance probes the quantum Mpemba effect when the inequality
$\text{Tr}({\rho^{\prime}}({t_{0}}){\texttt{R}_{1}})\leq\text{Tr}({\rho}({t_{0}}){\texttt{R}_{1}})\penalty 10000\ ,$ (B1)
is satisfied, where ${\rho^{\prime}}({t_{0}})=U\rho({t_{0}}){U^{\dagger}}$ is a quantum state obtained from the initial state $\rho({t_{0}})$ of the system, $U$ is a unitary matrix, and $\texttt{R}_{1}$} is the steady state of the system. Let $U={U_{1}}{P_{\pi}}{\Lambda^{\dagger}}$ a unitary matrix, where $P_{\pi}$} is an arbitrary permutation matrix, with ${P_{\pi}^{\dagger}}={P_{\pi}^{-1}}$. Here, $U_{1}$ and $\Lambda$ are unitary matrices formed by the eigenstates of the Hamiltonian $H$ and the probe state $\rho({t_{0}})$, respectively. The steady state for Davies map is given by the Gibbs state, ${\texttt{R}_{1}}=\exp(-\beta{H})/Z={U_{1}}\Sigma{U_{1}^{\dagger}}$, with $\Sigma=\exp(-\varepsilon/{k_{B}}T)/Z$, where $\varepsilon$ is the diagonal matrix formed by the eigenvalues of $H$, and $Z=\text{Tr}(\exp(-\varepsilon/{k_{B}}T))$ is the partition function. In this case, we have that
$\text{Tr}({\rho^{\prime}}({t_{0}}){\texttt{R}_{1}})$ $={Z^{-1}}\text{Tr}({U_{1}^{\dagger}}{U_{1}}{P_{\pi}}{\Lambda^{\dagger}}\Lambda{D}{\Lambda^{\dagger}}\Lambda{P_{\pi}^{\dagger}}{U_{1}^{\dagger}}{U_{1}}\Sigma)$
$={Z^{-1}}\text{Tr}({P_{\pi}}D{P_{\pi}^{\dagger}}\Sigma)\penalty 10000\ ,$ (B2)
where we have used the cyclic property of the trace, while
$\text{Tr}({\rho}({t_{0}}){\texttt{R}_{1}})$ $={Z^{-1}}\text{Tr}({U_{1}^{\dagger}}\Lambda{D}{\Lambda^{\dagger}}{U_{1}}\Sigma)$
$={Z^{-1}}\text{Tr}({U^{\prime}}{D}{U^{\prime\dagger}}\Sigma)\penalty 10000\ ,$ (B3)
where we define $U^{\prime}:={U_{1}^{\dagger}}\Lambda$. Hence, by substituting Eqs. (B) and (B) into Eq. (B1), we arrive at the inequality
$\text{Tr}({P_{\pi}}D{P_{\pi}^{\dagger}}\Sigma)\leq\text{Tr}({U^{\prime}}{D}{U^{\prime\dagger}}\Sigma)\penalty 10000\ .$ (B4)
To be clear, the role of the permutation matrix $P_{\pi}$} is to rearrange the entries along the diagonal of the matrix $D$. In detail, given $D=\text{diag}({\lambda_{1}},{\lambda_{2}},\ldots,{\lambda_{d}})$, we have that ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, where $\pi(j)=k$ defines the permutation operation mapping a given number $j$ into $k$, with $j,k\in\{1,\ldots,d\}$.
Next, we use that $\Sigma={\sum_{j}}\,{\alpha_{j}}|{j}\rangle\langle{j}|$ and $D={\sum_{j}}\,{\lambda_{j}}|{j}\rangle\langle{j}|$, where $\{|{j}\rangle\}_{j=1,\ldots,d}$} is the computational basis. In this case, the term $\text{Tr}({U^{\prime}}{D}{U^{\prime\dagger}}\Sigma)$ in the right-hand side of Eq. (B4) becomes
$\text{Tr}({U^{\prime}}{D}{U^{\prime\dagger}}\Sigma)={\sum_{j}}\,{\alpha_{j}}{\sum_{k}}\,{\lambda_{k}}{M_{jk}}\penalty 10000\ ,$ (B5)
where we define
${M_{jk}}:={|\langle{j}|{U^{\prime}}|{k}\rangle|^{2}}\penalty 10000\ .$ (B6)
We recognize ${M_{jk}}=\langle{j}|M|{k}\rangle$ as the $(j,k)$-th element of bistochastic matrix, since it fulfills the property ${\sum_{j}}\,{M_{jk}}={\sum_{k}}\,{M_{jk}}=1$. Indeed, by performing the sum over index $j$, we have that
${\sum_{j}}\,{M_{jk}}={\sum_{j}}\,{|\langle{j}|{U^{\prime}}|{k}\rangle|^{2}}={\sum_{j}}\,\langle{k}|{U^{\prime\dagger}}|{j}\rangle\langle{j}|{U^{\prime}}|{k}\rangle=1\penalty 10000\ ,$ (B7)
and also the sum over index $k$,
${\sum_{k}}\,{M_{jk}}={\sum_{k}}\,{|\langle{j}|{U^{\prime}}|{k}\rangle|^{2}}={\sum_{k}}\,\langle{j}|{U^{\prime}}|{k}\rangle\langle{k}|{U^{\prime\dagger}}|{j}\rangle=1\penalty 10000\ ,$ (B8)
where we have used that ${\sum_{j}}\,|{j}\rangle\langle{j}|=\mathbb{I}$. The same result holds for the sum over index $k$. In this case, according to the Birkhoff-von Neumann theorem, the matrix $M$ be recasted in terms of a convex combination of permutation matrices Ando (1989), i.e.,
$M={\sum_{l}}\,{\epsilon_{l}}{A_{l}}\penalty 10000\ ,$ (B9)
where $A_{l}$ is a given permutation matrix, with $0\leq{\epsilon_{l}}\leq 1$, and ${\sum_{l}}\,{\epsilon_{l}}=1$. In this case, one obtains
$\text{Tr}({U^{\prime}}{D}{U^{\prime\dagger}}\Sigma)$ $={\sum_{l}}\,{\epsilon_{l}}\,{\sum_{j,k}}\,{\alpha_{j}}{\lambda_{k}}\langle{j}|{A_{l}}|{k}\rangle$
$={\sum_{l}}\,{\epsilon_{l}}\,{\sum_{j,k}}\,{\alpha_{j}}{\lambda_{k}}{|\langle{j}|{A_{l}}|{k}\rangle|^{2}}\penalty 10000\ ,$ (B10)
where we used the fact that ${|\langle{j}|{A_{l}}|{k}\rangle|^{2}}=\langle{j}|{A_{l}}|{k}\rangle|$ since $A_{l}$} is a permutation matrix. It can be readily proved that
${\sum_{j,k}}\,{\alpha_{j}}{\lambda_{k}}{|\langle{j}|{A_{l}}|{k}\rangle|^{2}}$ $={\sum_{j,k}}\,{\alpha_{j}}{\lambda_{k}}\langle{j}|{A_{l}}|{k}\rangle\langle{k}|{A_{l}^{\dagger}}|{j}\rangle$
$=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)\penalty 10000\ .$ (B11)
Hence, by combining Eqs. (B) and (B), we obtain the result
$\text{Tr}({U^{\prime}}{D}{U^{\prime\dagger}}\Sigma)={\sum_{l}}\,{\epsilon_{l}}\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)\penalty 10000\ .$ (B12)
Finally, by combining Eqs. (B4) and (B12), one obtains
$\text{Tr}({P_{\pi}}{D}{P_{\pi}^{\dagger}}\Sigma)\leq{\sum_{l}}\,{\epsilon_{l}}\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)\penalty 10000\ .$ (B13)
The right-hand side of Eq. (B13) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{c_{l}}$ of the nonnegative elements ${c_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)$. We note that ${c_{\text{min}}}\leq{\sum_{l}}\,{\epsilon_{l}}{c_{l}}\leq{c_{\text{max}}}$, where ${c_{\text{min}}}=\text{min}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$ and ${c_{\text{max}}}=\text{max}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$, for all $0\leq{\epsilon_{l}}\leq 1$, with ${\sum_{l}}\,{\epsilon_{l}}=1$. This means that there exists two permutation matrices $P_{\pi,\text{max}}$} and $P_{\pi,\text{min}}$} that satisfy the upper bounds $\text{Tr}({P_{\pi,\text{max}}}{D}{P_{\pi,\text{max}}^{\dagger}}\Sigma)\leq{c_{\text{max}}}$ and $\text{Tr}({P_{\pi,\text{min}}}{D}{P_{\pi,\text{min}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$, respectively. The inequality in Eq. (B13) is satisfied in both scenarios, but we realize that the latter case provides a tighter upper bound than the former. In particular, there must exist an optimal permutation matrix $P_{\pi,\text{opt}}$} such that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$. To see this point, we note that
$\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)={\sum_{j=1}^{d}}\,{\alpha_{j}}{\lambda_{\pi(j)}}\penalty 10000\ ,$ (B14)
where $\pi(\bullet)$ maps index $j$ to $\pi(j)$, with $j,\pi(j)\in\{1,2,\ldots,d\}$. Therefore, to ensure that ${\sum_{j=1}^{d}}\,{\alpha_{j}}{\lambda_{\pi(j)}}\leq{c_{\text{min}}}$, it suffices to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\leq{\lambda_{\pi(2)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{n}}$ are listed in descending order, or vice-versa. Indeed, it is know from the so-called rearrangement inequality that ${\sum_{j=1}^{d}}\,{\alpha_{j}}{\lambda_{\pi(j)}}$ takes its minimum value if ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{d}}$ and ${\lambda_{\pi(1)}}\leq{\lambda_{\pi(2)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ Hardy et al. (1988). This proves the validity of the inequality in Eq. (B13), which also guarantees that Eq. (B1) is valid.
[…]